OpenStudy (anonymous):

A spherical snoowball is melting at a rate proportional to its surface area. That iis, the rate at which its volume is decreasing at any instant is proportional to its surface area at that rate. Prove that the radius of the snowball is decreasing at a constant rate.

OpenStudy (amistre64):

what formulas do you spose we should use for this?

OpenStudy (anonymous):

What formulas can you use?

OpenStudy (anonymous):

differentiation?

OpenStudy (amistre64):

my first thought is to try to recall the formulas for volume and surface area of a sphere

OpenStudy (amistre64):

differentiation is a method, which we will most likely use as well

OpenStudy (anonymous):

oksy, so consider volume of a sphere= 4/3 pi r^3 ?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

surface area

OpenStudy (amistre64):

good, \[V = \frac 43 \pi~r^3\] now we take the derivatve of that: \[V' = \frac {4(3)}{3} \pi~r^2~r'\]

OpenStudy (anonymous):

so we're not using pi r^2 ?

OpenStudy (anonymous):

why 43?

OpenStudy (anonymous):

oh sorry it's a fraction

OpenStudy (amistre64):

im thinking we will; now that we have the V' we are told that V' is equal to the formula for surface area (if im reading this right) 4(3) is just the result of taking an implicit derivative. the 3s cancels out and we are left with the formula for surface area, times r'

OpenStudy (amistre64):

\[V' = 4\pi r^2~r'\] we are told: \[V' = 4\pi~r^2\] so sub it in \[4\pi~r^2 = 4\pi r^2~r'\] now solve for r'

OpenStudy (amistre64):

is the latex coding showing up correctly on your side?

OpenStudy (anonymous):

no it's not but I think I'll figure it out.Thank you. :)

OpenStudy (amistre64):

V' = 4 pi r^2 r' we are told that V' = Surface area = 4 pi r^2, so sub it in 4 pi r^2 = 4 pi r^2 r' then solve for r'

OpenStudy (anonymous):

what is r'?

OpenStudy (anonymous):

the acceeration rate of the radius?

OpenStudy (dumbcow):

r' is dr/dt --> the rate the radius is changing over time acceleration is denoted as 2nd derivative